OpenStudy (misssmartiez):

Help me find the MAD of SOIL SAMPLE A: 61+ 61 + 62 + 65 +70 + 71 + 75 +81 +82 + 90 = 718 718/10 = 71.8 SOIL SAMPLE B 59 + 63 + 69 + 70 + 72 + 73 + 76 + 77 + 78 + 83 = 720 720/10= 72 (71.8 AND 72) IS THE MEAN,

OpenStudy (misssmartiez):

@RhondaSommer @rebeccaxhawaii @Directrix @math&ing001 Please, this is my last QUESTION!

OpenStudy (misssmartiez):

@mathmale @MrNood

OpenStudy (misssmartiez):

@Conqueror

OpenStudy (mathmale):

What could you possibly mean by MAD? Please explain.

OpenStudy (misssmartiez):

Mean of absolute deviation.

OpenStudy (misssmartiez):

I just need this, can you help me establish the MAD of both sample soil a and soil b, no one is HELPING me and I came up with myself, they just breeze over my questions.

OpenStudy (mathmale):

I assume you know what "absolute deviations" are. I'm not 100% sure on that. If yes, then you merely add up all the absolute deviations and divide your sum by the number of absolute deviations.

OpenStudy (misssmartiez):

No, I must have the mean of both soil A and B (Check B) ), subtract the mean from each value, and then find the mean of that, @mathmale, everyone is merely telling me things and not HELPING me through it, I have been STUCK for HOURS.

OpenStudy (conqueror):

Not good at math, Sorry.

OpenStudy (misssmartiez):

@mathmale A) To find the MAD, we must have the mean of both soil A and B (Check B) ), subtract the mean from each value, and then find the mean of that. 61 – 17.8 = 43.2 61 – 17.8 = 43.2 62 – 17.8 = 44.2 65 – 17.8 = 47.2 70 – 17.8= 52.2 71 – 17.8= 53.2 75 – 17.8= 57.2 81 – 17.8= 63.2 82 – 17.8= 64.2 90 – 17.8 = 72.2 43.2 x 2 + 44.2 + 47.2 + 52.2 + 53.2 + 57.2 + 63.2 + 64.2 + 72.2 = 540 540 / 10 = 54 MAD OF SOIL A: 54 I have so far

OpenStudy (mathmale):

Without having checked every detail of your work, I'd say you're on the right track.

OpenStudy (misssmartiez):

Thanks!